How would you make 300 mL of a 5.5 mM glucose solution?
Answer #1

Molecular weight of glucose = 180.156 g/mol

Concentration of required solution = 5.5 mM (this means 5.5 milimoles in 1 litre solution) = 5.5/1000 Molar  

Volume of solution required = 300 mL = 300/1000 Litre

Formula for the required mass of glucose for this solution:

Mass required = Molarity X Molecular Weight X Volume in Litre = (5.5/1000 X 180.156 X 300/1000) g

= 0.297 g

= 297 mg

Hence to make this solution, 297 mg of glucose is to be dissolved in 300 mL of distilled water.